# A queuing theory-based operating capacity model for multi-modal port operations

### v1

@@ Abstract
[2]organization=Coastal Hydraulics Laboratory, US Army Corps of Engineers Research and Development Center,addressline=3909 Halls Ferry Rd, city=Vicksburg, postcode=39180, state=MS, country=USA
@@ 1 Introduction
In these simulation and interaction model studies, maritime system capacity is typically defined as the maximum cargo or vessel throughput over a finite time period. Yet, this definition does not capture the idea of a long term, sustainable maximum throughput. In practice, ports may temporarily handle much higher vessel volumes per unit of time, but such conditions are unstable and cannot be sustained in the long run. This creates a need for a metric that defines the long-term operating capacity of multi-modal maritime transportation systems. Such a metric should account for the interactions between both waterside and landside systems and reflect the long-term freight-moving ability of the integrated port system. This metric would have value for stakeholders interested in understanding how infrastructural investments, such as channel expansion or terminal construction projects, may influence the overall capacity of an integrated port system. A metric capable of predicting the long-term outcomes of these initiatives and drawing comparisons between different project alternatives over an extended period would provide value to port planners and stakeholders.
@@
Contributions. The key contributions of this paper are as follows. We develop a queuing theory–based model for analyzing long-term port operations, which can be calibrated using real-world data. The model predicts operating capacity at both the port and terminal levels. It is broadly applicable to any port that handles one or more cargo types and has an anchorage area, followed by a channel leading to multiple terminals. Unlike many existing capacity models in the literature, our approach requires only data that can be readily obtained from archival Automatic Identification System (AIS) vessel trajectories and historical terminal logs, avoiding the need for extensive simulation or highly detailed datasets. The resulting capacity predictions can support stakeholders in disruption planning and in identifying the need for capacity expansion projects under future demand scenarios. To demonstrate the effectiveness and extensibility of the model, we present a detailed case study at the Port of Houston. In summary, our work makes three distinct contributions to the literature:
@@
    It presents a method for estimating long-term sustainable port operating capacity rather than maximum throughput, which is often unsustainable for multimodal ports over extended periods.
@@
    The proposed model can be calibrated using readily available archival AIS data and historical port data without requiring extensive data collection or simulation, and
@@
The remainder of the paper is structured as follows: Section 2 provides a brief review of relevant literature. Next, Section 3 presents the queuing models for each subsystem and introduces the framework for computing operating capacity. Section 4 compares the proposed techniques’ performance on empirical data for the Port of Houston. Finally, Section 5 summarizes our findings and suggests potential extensions to the research.
@@ 2 Literature review
Defining capacity in a multi-modal port system is challenging due to the complex interactions among multiple resources and parameters [8]. Consequently, there is no universally accepted definition of long-term port operating capacity [3]. Several studies have proposed capacity definitions for specific port components. Liu and Li [9] used arrangement theory to calculate the carrying capacity of navigable waterways by determining the number of ships that can be safely accommodated on the water. Fan and Cao [10] defined anchorage capacity as “the maximum number of vessels that can be accommodated by the anchorage over a period of time.” Using automatic identification system (AIS) vessel tracking data, Liu et al. [11] defined waterway capacity as “the ratio of the spatiotemporal resources of the port waterway to the spatiotemporal resources occupied by a single ship generally sailing in and out of the port waterway.” However, in real-world operations, vessels are rarely packed tightly, and waterway capacity is often constrained by factors beyond channel dimensions such as pilot and tug availability, downstream terminal resources, and navigation restrictions.
@@
Queuing models have been employed in a wide variety of domains to determine system capacities, including production [32], healthcare [33, 4], rail [34], and aerial vehicles [35]. As observed by Lantz and Rosén [7], capacity definitions typically stem from two types of approaches: engineering-based or time-based, both of which have inherent limitations. Engineering-based capacity measurements, such as estimating annual berth throughput from crane numbers, crane design rates, and assumed operating hours, are highly sensitive to input variables. Time-based studies in ports, on the other hand, may rely on measuring operational durations such as pilot boarding and vessel service times can introduce behavioral biases such as the Hawthorne effect [36]. These limitations are magnified when capacity is computed component-wise for an interdependent system. Queuing theory-based capacity computations [7] help mitigate these challenges while still considering all interdependencies. Zenzerović and Vilke [37] provided evidence that queuing models can be used to optimize capacities in container terminals, while Lee et al. [38] proposed a queuing model explicitly designed for the unique characteristics of container terminals to estimate capacity. However, a research gap exists in determining the long-term operating capacity of a multi-modal port system using queuing models and assessing the viability of such models in predicting the operating capacity of real-world systems. This research aims to fill this gap by proposing an operating capacity formulation by applying queuing models to the entire port system.
@@ 3 Port queuing model
To outline our multi-modal port queuing model, we first explain the life cycle of a vessel in a port and introduce the notations used in our model. We then propose queuing models that represent the waterway, terminal import, and terminal export processes. A vessel’s port life cycle begins with it entering the anchorage area and queuing until it can proceed to its destination terminal. In narrow waterways, a set of restrictions may govern navigation, such as combined beam, combined draft, and/or daylight. Once a berth is available and channel restrictions are met, the vessel requests resources (pilots and tugboats) to navigate to its terminal berth. Let $L_{a}^{i}$ denote the average number of ships waiting in the anchorage for vessel type $i$, where $i$ can be either container, break bulk, or liquid bulk vessels, and $W_{a}$ represent the average time vessels of all types spend waiting in the anchorage queue.
We employ a multiclass $M/M/1$ queuing model with equal priorities for the ship channel operation. Each class represents a cargo type, and since all vessels pass through the same channel, they are serviced at the same rate. However, each class of cargo arrives at a different arrival rate $\lambda_{a}^{i}$, following a Poisson process. We further assume that service times to enter the channel follow an exponential distribution with a mean service rate $\mu_{a}$, and departing vessels are prioritized over arriving vessels in the waterway channel. Consequently, we assume queues only form at the anchorage.
@@
$\lambda_{a}^{i}$ Vessel arrival rate to the system (at anchorage) for type $i$ vessels / hour
$L_{a}^{i}$ Mean length of queue of vessels of type $i$ at the anchorage vessels
@@
$W_{a}$ Mean waiting time at anchorage for all vessel types hour
@@ 3.1 The anchorage queuing model
We model the anchorage area using a multiclass $M/M/1$ queuing model, where the “server” represents an imaginary gate to the waterway channel. Each vessel requests service for channel entry, which is granted once all requirements, such as pilot and tug availability, navigation restrictions, and downstream terminal resources are satisfied. In this way, the service time distribution captures all channel restrictions and resource availability constraints that the vessels experience throughout the analysis period. The server is assumed to treat all vessels identically, with the same restrictions applying to every cargo class. This assumption allows us to compute a single effective service rate for the entire port, which we interpret as the port’s operating capacity. However, under this assumption, channel entry is indifferent to vessel class, so all vessel types experience the same average waiting time. Differences in expected queue lengths across classes arise solely from differences in their respective arrival rates. This is acceptable for our purpose, since we can estimate capacity based on queue lengths, and validation can still be performed using the mean waiting time across classes.
@@ 3.2 The terminal queuing model
The condition for the formed queue to be stable is given by Equation (4). The arrival rate $\lambda^{I}$ can be observed empirically from archival AIS data for a terminal. The import dwell time ($W_{c}^{I}$) and the moments of the batch size can be derived from historical port data. Equation (5) is then solved numerically to compute the cargo processing rate for imports subject to the stability constraint that $\rho_{c}^{I}<1$. The import process operating capacity is defined as the service rate of processing cargo that maintains the desired import cargo dwell over an extended period of stable queue formation.
@@
We employ a tri-tandem $M/M/S_{c}-M/M/1-M/M/1$ queuing model for export processes. Specifically, the gate processes are modeled as an $M/M/S_{c}$ system, while the holding area and yard processes are each modeled as an $M/M/1$ system. By applying Burke’s theorem [52], we assume cargo arrives at the terminal for export following a Poisson process. Empirical observations can provide the waiting times in the yard queue ($W_{c}^{E}$) and the cargo arrival rate at the gate $\lambda_{c}^{G}$. The service rate for exports can then be estimated using Equation (7) as derived by [7]. The export process operating capacity is defined as the service rate of processing cargo that maintains the desired export dwell time over an extended period of stable queue formation.
@@
Validation Wait time in anchorage ($W_{a}$) Yard Utilization ($L_{c}$) Queue length at gate* Queue length at holding* Yard utilization ($L_{c}$)
Table 2: Summary of queuing models used in this study. The table presents the specific queuing models, their application locations within the port, and the key inputs and outputs. The model inputs are categorized by quality of service (Q.O.S.) and arrival rates, and the model outputs are also shown. Validations marked with an asterisk $(^{*})$ are not performed in this study due to a lack of available data.
@@ 4 Results and discussions
To evaluate our proposed model for the port, we analyzed archival AIS data to determine input values for different cargo classes at the anchorage. The AIS-derived values for vessel arrival rates, mean anchorage wait times and anchorage queue lengths are shown in Table 3. The computed mean service rate was approximately $0.8$ vessels processed by the ship channel per hour. It is important to note that the mean service rate does not reflect the maximum number of vessels that the channel can accommodate at any given time. Instead, it indicates the mean inbound number of vessels that the port can handle over an extended period. For example, the computed $\mu_{a}$ values in Table 3 represent the average hourly processing rate of inbound vessels over each three-month period. The operating capacity value account for all operating conditions—including nights, weekends, holidays, channel restrictions, and terminal-level resource constraints. Thus, capacity values derived from this queuing model should only be compared against capacity values from the same queuing model for the same location under different conditions to assess various capacity scenarios.
@@
We observed that the port’s operating capacity remained relatively stable over the analysis period, although the mean anchorage waiting time varied from $32.97$ hours to $48.55$ hours, a variation of about $47\%$. Our model predicted the waiting time within close margins to the observed waiting times from archival AIS data, with relative errors ranging from $0.35\%$ in the fourth quarter of 2022 to $16.78\%$ in the fourth quarter of 2023. Table 3 summarizes the calculated service rate, traffic intensity, and anchorage queue length for the port. It is important to note that the traffic intensity during each analysis period was extremely high (close to 1), indicating that the entire port system was operating very near its capacity with minimal slack available. This, in turn, implies that even slight surges in arrivals could lead to disproportionately large increases in queue lengths.
@@
Excluding the analysis period from 2021 Q4 to 2022 Q3, the mean (standard deviation) operating capacity for the import process at BCT was 57.8 (4.4) containers per hour, while that for the export process was 74.6 (3.8) containers per hour. The results indicate that the export process had a 29% higher operating capacity than the import processes. It should be noted that the traffic intensity for the import process was approximately 0.88, whereas for the export process it approached 1. This indicates that the export operations are functioning very close to their effective capacity, while the import operations at BCT still retain considerable slack. Consequently, even slight surges in truck arrivals carrying export cargo could lead to significantly increased delays at the BCT terminal. These results are consistent with the observed higher dwell times at BCT, with the export process averaging about nine days compared to about four days for the import process [55].
@@ 5 Conclusion
We propose a queuing theory–based model to estimate operating capacities for multi-modal port processes without relying on data-intensive simulations. An advantage of queuing models is that they allow for modeling the entire port system rather than individual components. Furthermore, they provide a framework to incorporate port performance measures as quality-of-service variables in the capacity model. Unlike existing studies on port capacity, our model emphasizes long-term stable capacity—the throughput that can be sustained over an extended period of stable queue conditions—rather than the absolute maximum throughput a port can achieve under unstable conditions.
The approach requires only information on vessel arrival rates and queuing dynamics under operating conditions, which can be readily obtained from AIS data. Because the model relies on minimal and easily obtainable data, it is extendable to any port that has a structure similar to Figure 1 without the need to develop a detailed simulation. However, the model can predict capacities only under current resource availability, since it relies on real-world quality-of-service measures such as queue lengths and wait times. If a simulation is developed, these quality-of-service parameters could instead be derived from the simulation, allowing the model to predict operating capacity under different alternative scenarios.
@@
Nevertheless, our model has several limitations. First, several assumptions were made regarding the arrival process, service time distributions, and queuing disciplines. Although Poisson arrivals and exponentially distributed service times are common assumptions in queuing theory, several studies indicate that seaports may behave differently in practice. For instance, Legato and Mazza [61] noted that while Poisson processes and exponential service times are widely adopted, they may not accurately capture real-world port operations. Similarly, Demirci [47] observed that Erlang-distributed service times better reflect seaport behavior. In addition, our anchorage model assumes that all vessel classes are treated with equal priority and thus experience the same mean waiting time. This is an assumption that, while simplifying the analysis, may not hold in practice. Extending the capacity estimation techniques presented here to accommodate more general assumptions on arrivals, service distributions, and priority rules remains a promising direction for future research. Further, due to a lack of available data, we could not validate our models in greater detail on the terminal side. These models should be further tested with real data or a simulation to ensure their validity and extensibility. It should be emphasized that the goal of our work is not to propose the most realistic queuing model of port operations, but to demonstrate that queuing theory-based model provide a data-efficient framework for estimating port operating capacity.
@@
Second, queuing theory–based models omit several lower-level details, and much information is inevitably lost through aggregation. Although the presented models can predict port and terminal-level capacities, they do not account for important operational details such as inter-terminal vessel movements or the capacities of individual components. The model is instead based solely on aggregate quality-of-service parameters (such as dwell times and mean queue lengths). While such models have the advantage of requiring less data, this benefit can backfire if the underlying data are erroneous. This issue becomes particularly critical when terminals operate very close to, or effectively at, capacity. Under such conditions, queuing models, like the ones used in this study, become extremely sensitive to input data and small errors may translate into disproportionately large deviations in predicted performance. A promising direction for future work is to examine how uncertainty in input parameters affects capacity estimation, as vessel data are often subject to significant variability and measurement errors.
Acknowledgements
This work was partially funded through U.S. Army Engineer Research and Development Center (ERDC) contract #W912HZ23C0052, “Data-Driven, Multimodal Freight Modeling for Waterways and Ports.” The findings of this study reflect the authors’ work and are not official policies or standards of the U.S. Army Corps of Engineers. This work was also partially supported by the University Transportation Center (UTC) National Center for Understanding Future Travel Behavior and Demand (TBD).
The authors also thank Mr. Mark A. Cowan of ERDC for his valuable feedback throughout this work.
Author Contributions
The authors confirm their contribution to the paper as follows:
Conceptualization: D.B., K.B., S.B., K.M., M.A., M.K.
Methodology: D.B., K.B.
Data curation: D.B., K.B., M.K.
Formal analysis: D.B., K.B., S.B.
Writing—original draft: D.B., K.B.
Writing—review and editing: D.B., K.B., S.B., K.M., M.A., M.K.
All authors reviewed and approved the final version of the manuscript.

### v2

@@ Abstract
[2]organization=Coastal and Hydraulics Laboratory, US Army Corps of Engineers Research and Development Center,addressline=3909 Halls Ferry Rd, city=Vicksburg, postcode=39180, state=MS, country=USA
[…]
[3]organization=Applied Research Associates, Inc. for US Army Corps of Engineers Research and Development Center,addressline=3909 Halls Ferry Rd, city=Vicksburg, postcode=39180, state=MS, country=USA
@@ 1 Introduction
In these simulation and component-based studies, maritime system capacity is typically defined as the maximum cargo or vessel throughput over a finite time period. Yet, this definition does not capture the idea of a long term, sustainable maximum throughput. In practice, ports may temporarily handle much higher vessel volumes per unit of time, but such conditions are unstable and cannot be sustained in the long run. This creates a need for a metric that defines the long-term operating capacity of multi-modal maritime transportation systems. Such a metric should account for the interactions between both waterside and landside systems and reflect the long-term freight-moving ability of the integrated port system. This metric would have value for stakeholders interested in understanding how infrastructural investments, such as channel expansion or terminal construction projects, may influence the overall capacity of an integrated port system. A metric capable of predicting the long-term outcomes of these initiatives and drawing comparisons between different project alternatives over an extended period would provide value to port planners and stakeholders.
@@
Contributions. The key contributions of this paper are as follows. We develop a queuing theory–based model for analyzing long-term port operations, which can be calibrated using real-world data. The model predicts operating capacity at both the port and terminal levels. It is broadly applicable to any port that handles one or more cargo types and has an anchorage area, followed by a channel leading to multiple terminals. Unlike many existing capacity models in the literature, our approach requires only data that can be readily obtained from archival Automatic Identification System (AIS) vessel trajectories and historical terminal logs, avoiding the need for extensive simulation or highly detailed datasets. The resulting capacity predictions can support stakeholders in disruption planning and in identifying the need for capacity expansion projects under future demand scenarios. To demonstrate the effectiveness and extensibility of the model, we present a detailed case study at the Port of Houston. In summary, our work makes three main contributions to the literature, with the first two representing methodological advancements and the third demonstrating the practical value of our research:
@@
    It presents a method for estimating long-term sustainable port operating capacity rather than short-term maximum throughput, which is often unsustainable for multimodal ports over extended periods.
@@
    The proposed model can be calibrated using readily available archival AIS data and historical terminal logs without requiring extensive data collection or simulation, and
@@
The remainder of the paper is structured as follows: Section 2 provides a brief review of the relevant literature. Section 3 presents the queuing models for each subsystem and introduces the framework for computing operating capacity. Section 4 evaluates the performance of the proposed techniques using empirical data from the Port of Houston. Section 5 discusses practical applications of the model for port planning decisions. Finally, Section 6 summarizes our findings and outlines potential directions for future research.
@@ 2 Literature review
Defining capacity in a multi-modal port system is challenging due to the complex interactions among multiple resources and parameters [8]. Consequently, there is no universally accepted definition of long-term port operating capacity [3]. Several studies have proposed capacity definitions for specific port components. For instance, Liu and Li [9] used arrangement theory to calculate the carrying capacity of navigable waterways by determining the number of ships that can be safely accommodated on the water. Fan and Cao [10] defined anchorage capacity as “the maximum number of vessels that can be accommodated by the anchorage over a period of time.” Using automatic identification system (AIS) vessel tracking data, Liu et al. [11] defined waterway capacity as “the ratio of the spatiotemporal resources of the port waterway to the spatiotemporal resources occupied by a single ship generally sailing in and out of the port waterway.” However, in real-world operations, vessels are rarely packed tightly, and waterway capacity is often constrained by factors beyond channel dimensions such as pilot and tug availability, downstream terminal resources, and navigation restrictions.
@@
Queuing models have been employed in a wide variety of domains to determine system capacities, including production [32], healthcare [33, 4], rail [34], and aerial vehicles [35]. As observed by Lantz and Rosén [7], capacity definitions typically stem from two types of approaches: engineering-based or time-based, both of which have inherent limitations. Engineering-based capacity measurements, such as estimating annual berth throughput from crane numbers, crane design rates, and assumed operating hours, are highly sensitive to input variables. Time-based studies in ports, on the other hand, may rely on measuring operational durations such as pilot boarding and vessel service times can introduce behavioral biases such as the Hawthorne effect [36]. These limitations are magnified when capacity is computed component-wise for an interdependent system. Queuing theory-based capacity computations [7] help mitigate these challenges while still considering all interdependencies. Zenzerović and Vilke [37] provided evidence that queuing models can be used to optimize capacities in container terminals, while Lee et al. [38] proposed a queuing model explicitly designed for the unique characteristics of container terminals to estimate capacity. However, a research gap exists in determining the long-term operating capacity of a multi-modal port system using queuing models and assessing the viability of such models in predicting the operating capacity of real-world port systems. This research aims to fill this gap by proposing an operating capacity formulation by applying queuing models to the entire port system.
@@ 3 Port queuing model
To outline our multi-modal port queuing model, we first explain the life cycle of a vessel in a port and introduce the notations used in our model. We then propose queuing models that represent the waterway, terminal import, and terminal export processes. A vessel’s port life cycle begins with it entering the anchorage area and queuing until it can proceed to its destination terminal. In narrow waterways, a set of restrictions may govern navigation, such as combined beam, combined draft, and/or daylight. Once a berth is available and channel restrictions are met, the vessel requests resources (pilots and tugboats) to navigate to its terminal berth. Let $L_{a}^{i}$ denote the average number of ships waiting in the anchorage for vessel class $i$, where $i$ can be either container, break bulk, or liquid bulk vessels, and $W_{a}$ represent the average time vessels of all classes spend waiting in the anchorage queue.
We employ a multiclass $M/M/1$ queuing model with equal priorities for the ship channel operation. Each class represents a cargo type, and since all vessels pass through the same channel, they are serviced at the same rate. However, each class of cargo arrives at a different arrival rate $\lambda_{a}^{i}$, following a Poisson process. We further assume that service times to enter the channel follow an exponential distribution with a mean service rate $\mu_{a}$, and departing vessels are prioritized over arriving vessels in the waterway channel. Consequently, we assume vessel queues only form at the anchorage.
@@
$\lambda_{a}^{i}$ Vessel arrival rate to the system (at anchorage) for class $i$ vessels / hour
$L_{a}^{i}$ Mean length of queue of vessels of class $i$ at the anchorage vessels
@@
$W_{a}$ Mean waiting time at anchorage for all vessel classes hour
@@ 3.1 The anchorage queuing model
We model the anchorage area using a multiclass $M/M/1$ queuing model, where the “server” represents an imaginary gate to the waterway channel. Each vessel requests service for channel entry, which is granted once all requirements, such as pilot and tug availability, navigation restrictions, and downstream terminal resources are satisfied. In this way, the service time distribution captures all channel restrictions and resource availability constraints that the vessels experience throughout the analysis period. The server is assumed to treat all vessels identically, with the same restrictions applying to every cargo class. This assumption allows us to compute a single effective service rate for the entire port, which we interpret as the port’s operating capacity. However, under this assumption, channel entry is indifferent to vessel class, so all vessel classes experience the same average waiting time. Differences in expected queue lengths across classes arise solely from differences in their respective arrival rates. This is acceptable for our purpose, since we can estimate capacity based on queue lengths, and validation can still be performed using the mean waiting time across classes.
@@ 3.2 The terminal queuing model
The condition for the formed queue to be stable is given by Equation (4). The arrival rate of vessels at a terminal ($\lambda^{I}$) can be observed empirically from archival AIS data for a terminal. The import dwell time ($W_{c}^{I}$) and the moments of the batch size can be derived from historical port data. Equation (5) is then solved numerically to compute the cargo processing rate for imports subject to the stability constraint that $\rho_{c}^{I}<1$. The import process operating capacity is defined as the service rate of processing cargo that maintains the desired import cargo dwell over an extended period of stable queue formation.
@@
We employ a tri-tandem $M/M/S_{c}-M/M/1-M/M/1$ queuing model for export processes. Specifically, the gate processes are modeled as an $M/M/S_{c}$ system, while the holding area and yard processes are each modeled as an $M/M/1$ system. By applying Burke’s theorem [52] and assuming each truck brings only a unit of cargo, it follows that cargo arrives at the terminal for export following a Poisson process. Terminal logs can provide the export cargo waiting times in the yard ($W_{c}^{E}$) and the cargo arrival rate at the gate $\lambda_{c}^{G}$. The service rate for exports can then be estimated using Equation (7) as derived by [7]. The export process operating capacity is defined as the service rate of processing cargo that maintains the desired export dwell time over an extended period of stable queue formation.
@@
Validation Wait time in anchorage Yard utilization Queue length at gate* Queue length at holding* Yard utilization
Table 2: Summary of queuing models used in this study. The table presents the specific queuing models, their application locations within the port, the key inputs and outputs. The model inputs are categorized by quality-of-service (Q.O.S.) and arrival description. Validations marked with an asterisk $(^{*})$ are not performed in this study due to a lack of available data.
@@ 4 Results and discussions
To evaluate our proposed model for the port, we analyzed archival AIS data to determine input values for different cargo classes at the anchorage. The AIS-derived values for vessel arrival rates, mean anchorage wait times and anchorage queue lengths are shown in Table 3. The computed mean service rate was approximately $0.8$ vessels processed by the ship channel per hour. It is important to note that the mean service rate does not reflect the maximum number of vessels that the channel can accommodate at any given time. Instead, it indicates the mean inbound number of vessels that the port can handle over an extended period. For example, the computed $\mu_{a}$ values in Table 3 represent the maximum time-averaged hourly processing rate of inbound vessels over each three-month period. The operating capacity value account for all operating conditions—including nights, weekends, holidays, channel restrictions, and terminal-level resource constraints. Thus, capacity values derived from this queuing model should only be compared against capacity values from the same queuing model for the same location under different conditions to assess various capacity scenarios.
@@
We observed that the port’s operating capacity remained relatively stable over the analysis period, although the mean anchorage waiting time varied from $32.97$ hours to $48.55$ hours, a variation of about $47\%$. Our model predicted the waiting time within close margins to the observed waiting times from archival AIS data, with relative errors ranging from $0.35\%$ in the fourth quarter of 2022 to $16.78\%$ in the fourth quarter of 2023. Table 3 summarizes the calculated service rate, traffic intensity, and anchorage queue length for the port. It is important to note that the traffic intensity during each analysis period was extremely high (close to one), indicating that the entire port system was functioning near to its operating capacity with minimal slack available. This, in turn, implies that even slight surges in arrivals could lead to disproportionately large increases in queue lengths of vessels at the port anchorage.
@@
Excluding the analysis period from 2021 Q4 to 2022 Q3, the mean (standard deviation) operating capacity for the import process at BCT was 57.8 (4.4) containers per hour, while that for the export process was 74.6 (3.8) containers per hour. The results indicate that the export process had a 29% higher operating capacity than the import processes. Further, it should be noted that the traffic intensity for the import process was approximately 0.88, whereas for the export process it approached 1. This indicates that the export operations are functioning very close to its operating capacity, while the import operations at BCT still retain considerable slack. Consequently, even slight surges in truck arrivals carrying export cargo could lead to significantly increased delays at the BCT terminal. These results are consistent with the observed higher dwell times at BCT, with the export process averaging about nine days compared to about four days for the import process [55].
@@
5 Model application
The models presented in this study can be used to compute the long-term operating capacity for the entire port system (Section 3.1) and to compare the operating capacities of terminal import and export processes (Section 3.2). This section outlines how the model can be applied in practice to guide data-driven decisions on when and where to undertake capacity improvement projects.
Once the operating capacity of the entire port is determined, the presented queuing models can be used to predict how the port will perform under different vessel arrival rate pattern. For example, scenarios involving higher demand, surges in particular vessel classes, or shifts in the arrival distribution between vessel clasees can be evaluated using the anchorage model in Section 3.1. Model outputs such as traffic intensity, expected queue lengths, and wait times can then help determine whether capacity improvement projects should be considered. Similarly, terminals with the highest traffic intensities can be identified and prioritized for further analysis using the models in Section 3.2.
Because the model requires minimal data, it serves as an ideal initial screening tool to determine whether a detailed capacity assessment is needed and whether such an assessment should focus on the entire port or specific terminals. While making such decisions, it is important to note that ports often prefer to operate close to their operating capacity to maximize utilization of resources, as this directly translates into economic benefits. Consequently, relatively high traffic intensities and queue lengths, especially for cargo waiting in terminals, can often be tolerated. However, as traffic intensity approaches one, the system becomes increasingly sensitive to even slight changes in arrival patterns.
Once a decision has been made on which systems to evaluate further, a simulation can be developed to better understand system bottlenecks. Although we demonstrate the novelty of our approach within a simulation-free framework, using real-world quality-of-service measures to estimate capacity, it can also be used in conjunction with a simulations. For example, if the quality-of-service parameters are derived from a simulation rather than real-world data, the model can estimate operating capacity under different scenarios. In fact, using operating capacity alongside simulations, rather than relying solely on maximum throughput during a simulation horizon, provides a more realistic measure of sustainable performance. This supports better long-term infrastructure decisions by focusing on the port’s sustainable throughput rather than short-term peak loads that may not be maintainable.
@@ 6 Conclusion
We propose a queuing theory–based model to estimate operating capacities for multi-modal port processes without relying on data-intensive simulations. An advantage of queuing models is that they allow for modeling the entire port system rather than individual components. Furthermore, they provide a framework to incorporate port performance measures as quality-of-service variables in the capacity model. Unlike existing studies on port capacity, our model emphasizes long-term operating capacity—the throughput that can be sustained over an extended period of stable queue conditions—rather than the short-term maximum throughput a port can achieve under unstable conditions. The approach requires only information on vessel arrival rates and queuing dynamics under operating conditions, which can be readily obtained from AIS data and terminal logs. Because the model relies on minimal and easily obtainable data, it is extendable to any port with a structure similar to Figure 1 without the need for a detailed simulation.
@@
Nevertheless, our model has several limitations. First, several assumptions were made regarding the arrival process, service time distributions, and queuing disciplines. Although Poisson arrivals and exponentially distributed service times are common assumptions in queuing theory, several studies indicate that seaports may behave differently in practice. For instance, Legato and Mazza [61] noted that while Poisson processes and exponential service times are widely adopted, they may not accurately capture real-world port operations. Similarly, Demirci [47] observed that Erlang-distributed service times better reflect seaport behavior. In addition, our anchorage model assumes that all vessel classes are treated with equal priority and thus experience the same mean waiting time. This is an assumption that, while simplifying the analysis, may not hold in practice. Extending the capacity estimation techniques presented here to accommodate more general assumptions on arrivals, service distributions, and priority rules remains a promising direction for future research. Further, due to a lack of available data, we could not validate our models in greater detail on the terminal side. These models should be further tested with real data or a simulation to ensure their validity and extensibility. It should be emphasized that the goal of our work is not to propose the most realistic queuing model of port operations, but to demonstrate that queuing theory-based model provide a viable framework for estimating port operating capacity.
@@
Second, queuing theory–based models omit several lower-level details, and much information is inevitably lost through aggregation. Although the presented models can predict port and terminal-level capacities, they do not account for important operational details such as inter-terminal vessel movements or the capacities of individual components. The model is instead based solely on aggregate quality-of-service parameters (such as dwell times and mean queue lengths). While such models have the advantage of requiring less data, this benefit can backfire if the underlying data are erroneous. This issue becomes particularly critical when terminals operate very close to capacity. Under such conditions, queuing models, like the ones used in this study, become extremely sensitive to input data and small errors may translate into disproportionately large deviations in predicted performance. A promising direction for future work is to estimate port capacity with uncertain data, as vessel data are often subject to significant variability and measurement errors.